OpenStudy (anonymous):

Factor completely 2x^2 - 2x – 24. (2x - 6)(x + 4) (2x - 8)(x + 3) 2(x - 3)(x + 4) 2(x - 4)(x + 3)

OpenStudy (anonymous):

b?

OpenStudy (anonymous):

@skullpatrol

OpenStudy (anonymous):

d

OpenStudy (anonymous):

It can't be b, because (2x-8) can be factorized to 2(x-4)

OpenStudy (anonymous):

oh :/

OpenStudy (anonymous):

\(\large 2x^2 - 2x – 24=2(x^2-x-12)\)

OpenStudy (anonymous):

thanks guys c:

OpenStudy (anonymous):

\(\Large x^2-x-12=x^2+3x-4x-12=(x+3)(x-4)\)